# latexfontFont shape

### v1

@@ Abstract.
We define the class of explorable automata on finite or infinite words. This is a generalization of History-Deterministic (HD) automata, where this time non-deterministic choices can be resolved by building finitely many simultaneous runs instead of just one. We show that recognizing HD parity automata of fixed index among explorable ones is in PTime, thereby giving a strong link between the two notions. We then show that recognizing explorable automata is ExpTime-complete, in the case of finite words or parity automata up to index $[0,2]$. Additionally, we define the notion of $\omega$-explorable automata on infinite words, where countably many runs can be used to resolve the non-deterministic choices. We show ExpTime-completeness for $\omega$-explorability of automata on infinite words for the safety and coBüchi acceptance conditions. We finally characterize the expressivity of ($\omega$-)explorable automata with respect to the parity index hierarchy.
@@ 1. Introduction
One of the motivations to introduce the notion of explorability is to tackle one of the important open questions about HD automata: what is the complexity of deciding whether an automaton is HD? We explain in the following why explorability is relevant for this question, and show obstructions to some of our initial hopes in this direction.
Recognizing HD automata is known to be in PTime for Büchi [BK18] and coBüchi [KS15] automata, but even for $3$ parity ranks, the only known upper bound is ExpTime via the naive algorithm from [HP06]. We show how explorable automata can simplify this question: if the input automaton is explorable, then the problem becomes PTime for any fixed acceptance condition. Therefore, the question of recognizing HD automata can be shifted to: how hard is it to recognize explorable automata?
We then proceed to study the decidability and complexity of the explorability problem: deciding whether an input automaton on finite or infinite words is explorable. For this, we establish a connection with the population control problem studied in [BDG+19]. This problem asks, given an NFA with an arbitrary number of tokens in the initial state, whether a controller can choose input letters, thereby forcing every token to reach a designated state, even if tokens are controlled by an opponent. It is shown in [BDG+19] that the population control problem is ExpTime-complete, and we adapt their proof to our setting to show that the explorability problem is ExpTime-complete as well, already for NFAs. We also show that a direct reduction is possible, but at an exponential cost, yielding only a $2$-ExpTime algorithm for the NFA explorability problem. In the case of infinite words, we adapt the proof to the Büchi case, thereby showing that the Büchi explorability problem is in ExpTime as well. We also remark that, as in [BDG+19], the number of tokens needed to witness explorability can go as high as doubly exponential in the size of the automaton.
This ExpTime-completeness result means that we unfortunately cannot directly use the intermediate notion of explorable automata to improve on the complexity of recognizing HD automata in full generality, as could have been the hope. However, there can also be some frameworks where we can guarantee to obtain an explorable automaton, and therefore easily decide whether it is HD. A recent example of this from [BL22] is detailed in Section 2.5.2. More generally, we believe that this explorability notion is of interest towards a better understanding of non-determinism in automata theory.
@@
We also establish the expressivity of explorable automata, through reduction to or from deterministic parity automata. Surprisingly, the expressivity hierarchy of the explorable automata collapse, just as the hierarchy of non-determinist automata, albeit one step later. The general case is reached with parity-[1,3] explorable automata, which can recognize all regular languages.
@@
Summary of the contributions. We show that given an explorable parity automaton of fixed parity index, it is in PTime to solve its HDness problem, i.e. decide whether it is HD. The idea was already used in [BK18], and in [BL22] for quantitative automata. The algorithm used for Büchi HDness in [BK18] is conjectured to work for any acceptance condition (this is the “$G_{2}$ conjecture”), and it is in fact this algorithm that is shown here to work on any explorable parity automaton.
We show that given an NFA or a parity automaton with parities $\subseteq[0,2]$, it is decidable and ExpTime-complete to check whether it is explorable. We also study the expressivity in terms of recognized languages of the different parity classes of explorable automata. Our proof of ExpTime-completeness for NFA explorability uses techniques developed in [BDG+19], where ExpTime-completeness is shown for the NFA population control problem. We generalize this result to ExpTime explorability checking for parity [0,2] automata, requiring further adaptations. We also give a black box reduction using the result from [BDG+19]. This is enough to show decidability of the NFA explorability problem, but it yields a $2$-ExpTime algorithm. As in [BDG+19], the ExpTime algorithm yields a doubly exponential tight upper bound on the number of tokens needed to witness explorability. We show that deciding the explorability of parity [1,3] automata amounts to deciding the explorability of automata in the general parity case.
On infinite words, we show that any reachability automaton is $\omega$-explorable, but that this is not the case for safety automata. We show that both the safety and coBüchi $\omega$-explorability problems are ExpTime-complete. We also show that the Büchi case corresponds to the general case : any non-deterministic parity automaton can be converted in PTime to a Büchi automaton with same $\omega$-explorability status.
Related Works. Many works aim at quantifying the amount of non-determinism in automata. A survey by Colcombet [Col12] gives useful references on this question. Let us mention for instance the notion of ambiguity, which quantifies the number of simultaneous accepting runs. Similarly to [KM19], we can note that ambiguity is orthogonal to $k$-explorability. Remark however that our finite/countable/uncountable explorability hierarchy is reminiscent of the finite/polynomial/exponential ambiguity hierarchy (see e.g. [WS91]).
@@
The idea of $k$-explorability stems from the approach in [BK18], using games with tokens to tackle the HDness problem for Büchi automata. In this previous work, the idea of following a finite number of runs in parallel plays a central role in the proof. Remark however that the notion of explorability as studied here is stronger than what is needed in [BK18]. The $k$-explorability (and explorability) property was explicitly defined under the name $k$-History-Determinism in [BL22], as a proof tool to decide the HDness of LimInf and LimSup automata. The work [BL22] is part of a research effort to understand how partial determinism notions such as HDness play out in quantitative automata, see survey [Bok22]. Our goal here is to investigate explorability as defining a natural class of automata on finite and infinite words, somehow giving it an “official status” not restricted to an intermediate proof tool.
@@
The interest we took in the explorability notion originated in the fact that it makes deciding HDness much easier, and the hope was that by using this notion as an intermediate, we could obtain an algorithm improving on the ExpTime upper bound for deciding GFGness of parity automata of fixed index, e.g. to PSpace. As we described above, we ended up showing that this approach cannot yield an algorithm below ExpTime (at least not in full generality). However, although this was initially only a tool for this decision problem, explorability turns out to be a natural generalisation of HD automata, and an interesting class to study in itself. The first investigation of this notion, and in particular of its decidability, was the object of a short research internship by Milla Valnet under the supervision of the third author. It was expected that decidability of explorability would be a reachable goal for such a short internship, but it turned out that this was overly optimistic. The internship yielded preliminary results, and in particular was useful to introduce and study the notion of “coverability”. This version of the problem does not take acceptance conditions into account, but only asks that at any point of the run, every state that could be reached is actually occupied by a token. After the internship, we continued to use this coverability notion as a stepping stone towards an understanding of explorability. However, after more preliminary results and unsuccessful attempts at obtaining decidability, we discovered the connection between explorability and population control from [BDG+19], that rendered the intermediate notion of coverability useless for our purposes, and we then focused on exploiting that link. We chose to leave coverability out of the present exposition, as it feels like a “watered-down” version of explorability, but it could be useful in some contexts, hence we briefly mention it in this chronological account. Let us just informally state here that it is straightforward to modify our proofs in order to show that deciding whether an NFA is coverable is ExpTime-complete as well. The first results were obtained during the PhD of Emile Hazard, and published in [HK23]. Some new results, namely ExpTime algorithm for coBüchi and $[0,2]$-explorability, and expressivity results, were obtained during the internship of Olivier Idir. This paper, extending [HK23], aims at gathering what we currently know about explorable automata.
@@ 2.2. Automata
[…]
@@
Transitions from $F$ will be called Büchi transitions in Büchi automata, and transition from $\Delta\setminus F$ will be called coBüchi transitions in coBüchi automata.
@@
For all these acceptance conditions on infinite words, we will sometimes use state-based acceptance instead of transition-based when more convenient for our constructions. Recall that we can switch from transition-based to state-based with a doubling of the number of states, and the translation from state-based to transition-based can be done without changing the size of the automaton. These translations do not affect any of the explorability properties considered in this paper. See [Cas23] for details on the merits of transition-based acceptance conditions over the state-based ones.
@@ 2.3. Games
A game $\mathcal{G}=(V_{0},V_{1},v_{I},E,W)$ of infinite duration between two players 00 and $1$ consists of: a finite set of positions $V$ being a disjoint union of $V_{0}$ and $V_{1}$; an initial position $v_{I}\in V$; a set of edges $E\subseteq V\times V$; and a winning condition $W\subseteq V^{\omega}$. We will later use names more explicit than 00 and $1$ for the players, describing their roles in the various games we will define.
A play is an infinite sequence of positions $v_{0}v_{1}v_{2}\dots\in V^{\omega}$ such that $v_{0}=v_{I}$ and for all $n\in\mathbb{N}$, $(v_{n},v_{n+1})\in E$. A play $\pi\in V^{\omega}$ is winning for Player 0 if it belongs to $W$. Otherwise $\pi$ is winning for Player 1.
A strategy for Player 00 (resp. $1$) is a function ${\sigma_{0}}:{V^{\ast}\times V_{0}}\to{V}$ (resp. ${\sigma_{1}}:{V^{\ast}\times V_{1}}\to{V}$), describing which edge should be played given the history of the play $u\in V^{\ast}$ and the current position $v\in V$. A strategy $\sigma_{P}$ has to obey the edge relation, i.e. there has to be an edge in $E$ from $v$ to $\sigma_{P}(u,v)$. A play $\pi=v_{0}v_{1}v_{2}\dots$ is consistent with a strategy $\sigma_{P}$ of a player $P$ if for every $n$ such that $v_{n}\in V_{P}$ we have $v_{n+1}=\sigma_{P}(v_{0}\ldots v_{n-1},v_{n})$.
A strategy for Player 00 (resp. Player $1$) is positional (or memoryless) if it does not use the history of the play, i.e. it can be seen as a function $V_{0}\to V$ (resp. $V_{1}\to V$).
@@
In the interest of readability, when describing games in the paper, we will not give explicit definitions of the sets $V_{0}$, $V_{1}$ and $E$, but give slightly more informal descriptions in terms of possible actions of players at each round. It is straightforward to build a formal description of the games from such a description.
@@ 2.4. Explorability
[$k$-explorability game] Consider a non-deterministic automaton $\mathcal{A}$ on finite or infinite words, and an integer $k$. The $k$-explorability game on $\mathcal{A}$ is played on the arena $Q^{k}$. The two players are called Determiniser and Spoiler, and they play as follows.
@@
Consider the automata from Figure 1. The NFA $\mathcal{A}_{k}$ on alphabet $\{a,a_{1},\dots,a_{k}\}$ is $k$-explorable, but not $(k-1)$-explorable. It can easily be adapted to a binary alphabet, by replacing in the automaton $a_{1},\dots,a_{k}$ by distinct words of the same length.
On the other hand, the NFA $\mathcal{C}$ is a non-explorable NFA accepting all words on alphabet $\Sigma=\{a,b\}$. Indeed, Spoiler can win the $k$-explorability game for all $k$, by eliminating tokens one by one, choosing at each step the letter $b$ if $q_{1}$ is occupied by at least one token, and the letter $a$ otherwise.
@@
The NFA $\mathcal{B}_{k}$ from Figure 2 with $13k+$ 1 states on alphabet $\Sigma=\{a,b\}$ is explorable, but requires $2^{k}$ tokens. Indeed, since when choosing the $2i^{\mathrm{th}}$ letter Spoiler can always pick the state $p_{i}$ or $r_{i}$ containing the least amount of tokens to decide whether to play $a$ or $b$, the best strategy for Determiniser is to split his tokens evenly at each $q_{i}$. This means he needs to start with $2^{k}$ tokens to end up with at least one token in $q_{k}$ after a word of $\Sigma^{2k}$.
@@
  • Any automaton for a finite language is explorable.
@@
If $\mathcal{A}$ is $k$-explorable and $n\geq k$, then Determiniser can win the $n$-explorability game by using the same strategy with the first $k$ tokens and making arbitrary choices with the $n-k$ remaining tokens.
@@
However, Determiniser may need up to $kn$ tokens to play in $\mathcal{A}\times\mathcal{B}$ when the accepting set is $F_{\mathcal{A}}\times F_{\mathcal{B}}$: the token $(i,j)$ will use the strategy of the token $i$ in the $k$-explorability game of $\mathcal{A}$ together with the strategy of the token $j$ in the $n$-explorability game of $\mathcal{B}$. This lower bound of $kn$ cannot be improved: consider for instance the intersection product $\mathcal{A}_{k}\times\mathcal{A}_{n}$, where $\mathcal{A}_{k},\mathcal{A}_{n}$ are from Example 2.4, using as alphabet the cartesian product of their respective alphabets: $\{a,a_{1},a_{2},\dots,a_{k}\}\times\{a,a_{1},\dots,a_{n}\}$, or $\{a,b\}^{2}$ in their binary alphabet versions.
@@
Given an NFA $\mathcal{A}$ and an integer $k$, it is ExpTime-complete to decide whether $\mathcal{A}$ is $k$-explorable (even if we fix $k=|Q_{\mathcal{A}}|/2$).

### v2

# Explorable Parity Automata

@@ Abstract.
We define the class of explorable automata on finite or infinite words. This is a generalisation of History-Deterministic (HD) automata, where this time non-deterministic choices can be resolved by building finitely many simultaneous runs instead of just one. We show that recognizing HD parity automata of fixed index among explorable ones is in PTime, thereby giving a strong link between the two notions. We then show that recognizing explorable automata is ExpTime-complete, in the case of finite words or parity automata up to index $[0,2]$. Additionally, we define the notion of $\omega$-explorable automata on infinite words, where countably many runs can be used to resolve the non-deterministic choices. We show ExpTime-completeness for $\omega$-explorability of automata on infinite words for the safety and coBüchi acceptance conditions. We finally characterize the expressivity of ($\omega$-)explorable automata with respect to the parity index hierarchy.
@@ 1. Introduction
We then proceed to study the decidability and complexity of the explorability problem: deciding whether an input automaton on finite or infinite words is explorable. For this, we establish a connection with the population control problem studied in [BDG+19]. This problem asks, given an NFA with an arbitrary number of tokens in the initial state, whether a controller can choose input letters, thereby forcing every token to reach a designated state, even if tokens are controlled by an opponent. It is shown in [BDG+19] that the population control problem is ExpTime-complete, and we adapt their proof to our setting to show that the explorability problem is ExpTime-complete as well, already for NFAs. We also show that a direct reduction is possible, but at an exponential cost, yielding only a $2$-ExpTime algorithm for the NFA explorability problem. In the case of infinite words, we adapt the proof to $[0,2]$-parity case, thereby showing that the $[0,2]$-explorability problem is in ExpTime as well. We also remark that, as in [BDG+19], the number of tokens needed to witness explorability can go as high as doubly exponential in the size of the automaton.
@@
We also establish the expressivity of explorable automata, through reduction to or from deterministic parity automata. Surprisingly, the expressivity hierarchy of the explorable automata collapse, just as the hierarchy of non-deterministic automata, albeit one step later. The general case is reached with $[1,3]$-parity explorable automata, which can recognize all regular languages.
@@
Summary of the contributions.
We show that given an NFA or a parity automaton with parities $\subseteq[0,2]$, it is decidable and ExpTime-complete to check whether it is explorable. We also study the expressivity in terms of recognized languages of the different parity classes of explorable automata. Our proof of ExpTime-completeness for NFA explorability uses techniques developed in [BDG+19], where ExpTime-completeness is shown for the NFA population control problem. We generalise this result to ExpTime explorability checking for $[0,2]$-parity automata, requiring further adaptations. We also give a black box reduction using the result from [BDG+19]. This is enough to show decidability of the NFA explorability problem, but it yields a $2$-ExpTime algorithm. As in [BDG+19], the ExpTime algorithm yields a doubly exponential tight upper bound on the number of tokens needed to witness explorability. We show that deciding the explorability of $[1,3]$-parity automata amounts to deciding the explorability of automata in the general parity case.
On infinite words, we show that any reachability automaton is $\omega$-explorable, but that this is not the case for safety automata. We show that both the safety and coBüchi $\omega$-explorability problems are ExpTime-complete. We also show that the Büchi case corresponds to the general case : any nondeterministic parity automaton can be converted in PTime to a Büchi automaton with same $\omega$-explorability status.
Related Works.
Zooming out, many works aim at quantifying the amount of non-determinism in automata. A survey by Colcombet [Col12] gives useful references on this question. Let us mention for instance the notion of ambiguity, which quantifies the number of simultaneous accepting runs. Similarly to [KM19], we can note that ambiguity is orthogonal to $k$-explorability. Remark however that our finite/countable/uncountable explorability hierarchy is reminiscent of the finite/polynomial/exponential ambiguity hierarchy (see e.g. [WS91]).
@@
The particular problem of deciding in polynomial time whether a given automaton is HD has attracted a lot of attention. After some initial results on NFA, Büchi and coBüchi conditions [LR13, KS15, BK18, BKLS20], the general result was recently obtained by Lehtinen and Prakash in [LP25], showing that for any fixed parity index (and consequently for any fixed $\omega$-regular condition) it is decidable in polynomial time whether a given parity automaton is HD. It is shown in [LP25] that the two-token game introduced in [BK18] for Büchi automata actually characterizes HDness for any acceptance condition, as conjectured in [BK18]. The inspiration for $k$-explorability stems from this kind of approach, where the idea of following a finite number of runs in parallel plays a central role. Remark however that the notion of explorability as studied here is stronger than what is needed in the two-token game from [BK18, LP25]. The $k$-explorability (and explorability) property was explicitly defined under the name $k$-History-Determinism in [BL22], as a proof tool to decide the HDness of LimInf and LimSup automata. The work [BL22] is part of a research effort to understand how partial determinism notions such as HDness play out in quantitative automata, see survey [Bok22]. Our goal here is to investigate explorability as defining a natural class of automata on finite and infinite words, somehow giving it an “official status” not restricted to an intermediate proof tool.
@@
The interest we took in the explorability notion originated in the fact that it makes deciding HDness much easier, and the hope was that by using this notion as an intermediate, we could obtain an algorithm improving on the ExpTime upper bound for deciding HDness of parity automata of fixed index, e.g. to PSpace. As we described above, we ended up showing that this approach cannot yield an algorithm below ExpTime (at least not in full generality). However, although this was initially only a tool for this decision problem, explorability turns out to be a natural generalisation of HD automata, and an interesting class to study in itself. The first investigation of this notion, and in particular of its decidability, was the object of a short research internship by Milla Valnet under the supervision of the third author. It was expected that decidability of explorability would be a reachable goal for such a short internship, but it turned out that this was overly optimistic. The internship yielded preliminary results, and in particular was useful to introduce and study the notion of “coverability”. This version of the problem does not take acceptance conditions into account, but only asks that at any point of the run, every state that could be reached is actually occupied by a token. After the internship, we continued to use this coverability notion as a stepping stone towards an understanding of explorability. However, after more preliminary results and unsuccessful attempts at obtaining decidability, we discovered the connection between explorability and population control from [BDG+19], that rendered the intermediate notion of coverability useless for our purposes, and we then focused on exploiting that link. We chose to leave coverability out of the present exposition, as it feels like a “watered-down” version of explorability, but it could be useful in some contexts, hence we briefly mention it in this chronological account. Let us just informally state here that it is straightforward to modify our proofs in order to show that deciding whether an NFA is coverable is ExpTime-complete as well. The first results were obtained during the PhD of Emile Hazard, and published in [HK23]. Some new results, namely ExpTime algorithm for coBüchi and $[0,2]$-explorability, and expressivity results, were obtained during the internship of Olivier Idir. This paper, extending [HK23], aims at gathering what we currently know about explorable automata.
@@ 2.2. Automata
We work with a fixed finite alphabet $\Sigma$. We will use the following default notation for the components of an automaton $\mathcal{A}$: $Q_{\mathcal{A}}$ for its set of states, $q_{0}^{\mathcal{A}}$ for its initial state, $F_{\mathcal{A}}$ for its accepting states, $\Delta_{\mathcal{A}}$ for its set of transitions. If the automaton is clear from context, the subscript/superscript $\mathcal{A}$ might be omitted. We might also specify its alphabet by $\Sigma_{\mathcal{A}}$ instead of $\Sigma$ for cases where different alphabets come into play. If $\Delta\subseteq Q\times\Sigma\times Q$ is the transition relation, and $(p,a)\in Q\times\Sigma$, we will note $\Delta(p,a)=\{q\in Q,(p,a,q)\in\Delta\}$. If $X\subseteq Q$, we note $\Delta(X,a)=\bigcup_{p\in X}\Delta(p,a)$. A transition $(p,a,q)$ will often be noted $p\overset{a}{\longrightarrow}q$.
@@
Transitions from $F$ will be called Büchi transitions in Büchi automata, and transitions from $\Delta\setminus F$ will be called coBüchi transitions in coBüchi automata.
@@
For all these acceptance conditions on infinite words, we will sometimes use state-based acceptance instead of transition-based when more convenient for our constructions. Recall that we can switch from transition-based to state-based via a state blow-up by a factor of number of priorities (e.g. doubling the number of states for Büchi or coBüchi condition), and the translation from state-based to transition-based can be done without changing the size of the automaton. These translations do not affect any of the explorability properties considered in this paper. See [Cas23] for details on the merits of transition-based acceptance conditions over the state-based ones.
@@ 2.3. Games
A game $\mathcal{G}=(V_{0},V_{1},v_{I},E,W)$ of infinite duration between two players 0 and $1$ consists of: a finite set of positions $V$ being a disjoint union of $V_{0}$ and $V_{1}$; an initial position $v_{I}\in V$; a set of edges $E\subseteq V\times V$; and a winning condition $W\subseteq V^{\omega}$. We will later use names more explicit than 0 and $1$ for the players, describing their roles in the various games we will define.
A play is an infinite sequence of positions $v_{0}v_{1}v_{2}\dots\in V^{\omega}$ such that $v_{0}=v_{I}$ and for all $n\in\mathbb{N}$, $(v_{n},v_{n+1})\in E$. A play $\pi\in V^{\omega}$ is winning for Player 0 if it belongs to $W$. Otherwise, $\pi$ is winning for Player 1.
A strategy for Player 0 (resp. $1$) is a function ${\sigma_{0}}:{V^{\ast}\times V_{0}}\to{V}$ (resp. ${\sigma_{1}}:{V^{\ast}\times V_{1}}\to{V}$), describing which edge should be played given the history of the play $u\in V^{\ast}$ and the current position $v\in V$. A strategy $\sigma_{P}$ has to obey the edge relation, i.e. there has to be an edge in $E$ from $v$ to $\sigma_{P}(u,v)$. A play $\pi=v_{0}v_{1}v_{2}\dots$ is consistent with a strategy $\sigma_{P}$ of a player $P$ if for every $n$ such that $v_{n}\in V_{P}$ we have $v_{n+1}=\sigma_{P}(v_{0}\ldots v_{n-1},v_{n})$.
A strategy for Player 0 (resp. Player $1$) is positional (or memoryless) if it does not use the history of the play, i.e. it can be seen as a function $V_{0}\to V$ (resp. $V_{1}\to V$).
@@
We could equivalently define a winning condition on edges: $W\subseteq E^{\omega}$, but we will not need to dive into this distinction here. This is because in the interest of readability, when describing games in the paper, we will not give explicit definitions of the sets $V_{0}$, $V_{1}$ and $E$, but give slightly more informal descriptions in terms of possible actions of players at each round. It is straightforward to build a formal description of the games from such a description.
@@ 2.4. Explorability
[$k$-explorability game] Consider a non-deterministic automaton $\mathcal{A}$ on finite or infinite words, and an integer $k$. The $k$-explorability game on $\mathcal{A}$ is played on the arena $Q^{k}$. We will consider that a position $S=(q_{1},\dots,q_{k})$ of the arena $Q^{k}$ is described via $k$ tokens, numbered $1$ to $k$, and placed respectively in the states $q_{1},\dots,q_{k}$ of $\mathcal{A}$. If $l\in[1,k]$ is the index of a token, we will note $S(l)=q_{l}$ the location of this token in position $S$. The two players are called Determiniser and Spoiler, and they play as follows.
@@
Consider the automata from Figure 1. The NFA $\mathcal{A}_{k}$ on alphabet $\{a,a_{1},\dots,a_{k}\}$ is $k$-explorable, but not (k1)(k-1)-explorable. It can easily be strengthened to a binary alphabet whilst still requiring $k$ tokens, by replacing the transition labels $a_{1},\dots,a_{k}$ by distinct words of the same length.
On the other hand, the NFA $\mathcal{C}$ is a non-explorable NFA accepting all words on alphabet $\Sigma=\{a,b\}$. Indeed, Spoiler can win the $k$-explorability game for all $k$, by sending tokens to the sink $q_{3}$ one by one, choosing at each step the letter $b$ if $q_{1}$ is occupied by at least one token, and the letter $a$ otherwise.
@@
The NFA $\mathcal{B}_{k}$ from Figure 2 with $3k+1$ states on alphabet $\Sigma=\{a,b\}$ is explorable, but requires $2^{k}$ tokens. Indeed, since when choosing the $2i^{\mathrm{th}}$ letter Spoiler can always pick the state $p_{i}$ or $r_{i}$ containing the least amount of tokens to decide whether to play $a$ or $b$, the best strategy for Determiniser is to split his tokens evenly at each $q_{i}$. This means he needs to start with $2^{k}$ tokens to end up with at least one token in $q_{k}$ after a word of $\Sigma^{2k}$.
@@
  • Any automaton for a finite language (i.e. containing finitely many words) is explorable.
@@
If $\mathcal{A}$ is $k$-explorable and $n\geq k$, then Determiniser can win the $n$-explorability game by using the same strategy with the first $k$ tokens and making arbitrary choices with the nkn-k remaining tokens.
@@
However, Determiniser may need up to $kn$ tokens to play in $\mathcal{A}\times\mathcal{B}$ when the accepting set is $F_{\mathcal{A}}\times F_{\mathcal{B}}$, as one token needs to accept both paths of the product. The token $(i,j)$ will use the strategy of the token $i$ in the $k$-explorability game of $\mathcal{A}$ together with the strategy of the token $j$ in the $n$-explorability game of $\mathcal{B}$. This lower bound of $kn$ cannot be improved: consider for instance the intersection product $\mathcal{A}_{k}\times\mathcal{A}_{n}$, where $\mathcal{A}_{k},\mathcal{A}_{n}$ are from Example 2.4, using as alphabet the cartesian product of their respective alphabets: $\{a,a_{1},a_{2},\dots,a_{k}\}\times\{a,a_{1},\dots,a_{n}\}$, or $\{a,b\}^{2}$ in their binary alphabet versions.
@@
Given an NFA $\mathcal{A}$ and an integer $k$ given in unary, it is ExpTime-complete to decide whether $\mathcal{A}$ is $k$-explorable (even if we fix $k=|Q_{\mathcal{A}}|/2$).
The upper bound follows from the fact that the $k$-explorability game of an NFA is a safety game of size $|\mathcal{A}|^{k}\times|\mathcal{D}|$, where $\mathcal{D}$ is a determinization of $\mathcal{A}$. Overall, this game is exponential in $|\mathcal{A}|+k$. The lower bound is obtained by reducing from an ExpTime-complete combinatorial game from [SC79], which generalizes SAT by allowing two players to alternately choose values for variables. This game is encoded into the $k$-explorability game of an NFA $\mathcal{A}$, with $k$ being half the number of states of $\mathcal{A}$.